# Disen Lan    Weigao Sun    Jiaxi Hu    Jusen Du    Yu Cheng

### v1

@@ Abstract
Transformers with linear recurrent modeling offer linear-time training and constant-memory inference. Despite their demonstrated efficiency and performance, pretraining such non-standard architectures from scratch remains costly and risky. The linearization of large language models (LLMs) transforms pretrained standard models into linear recurrent structures, enabling more efficient deployment. However, current linearization methods typically introduce additional feature map modules that require extensive fine-tuning and overlook the gating mechanisms used in state-of-the-art linear recurrent models. To address these issues, this paper presents Liger, short for Linearizing LLMs to gated recurrent structures. Liger is a novel approach for converting pretrained LLMs into gated linear recurrent models without adding extra parameters. It repurposes the pretrained key matrix weights to construct diverse gating mechanisms, facilitating the formation of various gated recurrent structures while avoiding the need to train additional components from scratch. Using lightweight fine-tuning with Low-Rank Adaptation (LoRA), Liger restores the performance of the linearized gated recurrent models to match that of the original LLMs. Additionally, we introduce Liger Attention, an intra-layer hybrid attention mechanism, which significantly recovers 93% of the Transformer-based LLM at 0.02% pre-training tokens during the linearization process, achieving competitive results across multiple benchmarks, as validated on models ranging from 1B to 8B parameters. Code is available at https://github.com/OpenSparseLLMs/Linearization.
@@ 1 Introduction
Large language models (LLMs) have demonstrated exceptional performance across various natural language processing tasks (Chintala, 2023; Team, 2023; Zhu et al., 2024; Qu et al., 2024). However, the Transformer-based architecture (Vaswani et al., 2017) used in modern LLMs, with its reliance on softmax attention, suffers from quadratic computational complexity. This inefficiency results in significant speed and memory challenges, particularly during pretraining on long sequences. During inference, the Key-Value (KV) cache (Kwon et al., 2023) grows linearly with the input sequence length, leading to reduced inference speed and high memory usage, which severely limits the capability of these models for handling long-sequence tasks (Sun et al., 2024b). In contrast, models based on linear recurrent modeling (Katharopoulos et al., 2020; Yang et al., 2023; Qin et al., 2024b; Sun et al., 2024a; Du et al., 2025) provide linear-time training and constant-memory inference, offering substantial efficiency benefits and positioning themselves as promising candidates for the next generation of foundational architectures (MiniMax et al., 2025).
@@
While pretraining LLMs using architectures based on linear resurrent modeling reduces costs due to their linear training complexity, the high expenses of pretraining from scratch associated with large model sizes and datasets still remain a major obstacle to their adoption and practical use. This challenge has hindered the advancement of linear recurrent models. Linearizing pretrained LLMs like SUPRA (Mercat et al., 2024), MambaInLlama (Wang et al., 2024) and LoLCATs (Zhang et al., 2024a), as an emerging new direction, allows the transfer of weights from an existing pretrained model to one with linear recurrent modeling architectures at a small fraction of the original pretraining cost. The linearization approach is a promising post-training technique to enable efficient pretrained model deployment while preserving their performance. Gating mechanisms (Qin et al., 2024a; Sun et al., 2023) play a crucial role in linear recurrent models by controlling memory retention and forgetting, with their effectiveness widely demonstrated in such architectures. However, incorporating gate modules as additional components requires both transferring weights from pre-trained LLMs and training these gating modules from scratch. This process not only increases the cost of linearization but also creates a larger architectural divergence from Transformer-based LLMs. This divergence may hinder the effective approximation of softmax attention, limiting the performance of gated linear recurrent models (Zhang et al., 2024d). Moreover, existing linearization methods often overlook the detailed design considerations of gated linear models, and the newly added modules fail to leverage the pre-trained weights of LLMs, further reducing the efficiency of linearization.
@@
In this paper, we present Liger, which stands for Linearizing large language models into gated recurrent structures, a novel approach for linearizing LLMs. Liger repurposes the weights from pre-trained Transformer-based LLMs and introduces a novel method for constructing crucial gating mechanisms in gated recurrent structures using the key projection. This approach avoids the complex attention transfer process found in existing linearization methods. After transforming the weights and constructing the gating mechanisms, Liger requires only lightweight fine-tuning of the linearized gated recurrent model parameters through LoRA autoregressive training. This efficient process restores the model’s performance with minimal linearization cost, achieving competitive results across a range of language modeling and understanding benchmarks while benefiting from the linear-time inference efficiency of the recurrent architecture.
@@
  • We apply Liger to linearize the latest Llama3 series, ranging from 1B to 8B parameters. Experimental results show that Liger outperforms existing linearization methods (like SUPRA (Mercat et al., 2024), MambaInLlama (Wang et al., 2024) and LoLCATs (Zhang et al., 2024a)), in terms of both efficiency and its ability to preserve the original performance of pre-trained LLMs.
@@ 2 Preliminary
The standard softmax attention is highly rely on the growing KV Cache (Chou et al., 2024; Wang et al., 2024) to recall the history "memory" for sequence modeling, which results in quadratic complexity and costly memory requirements especially in long context setting.
@@
Although linear attention with causal mask matrix cannot use matrix associativity to reduce the parallel form training complexity from quadratic to linear, its chunk-wise parallel form allows hardware-efficient sub-quadratic and partially parallel training (Yang et al., 2023; Sun et al., 2024b, a; Qin et al., 2024a).
@@ 3.1 Gate Construction by Key Projection
Gated Retention (Sun et al., 2024c) $\mathbf{G}_{t}=\alpha_{t}\mathbf{1}^{\top}\mathbf{1}$ $\alpha_{t}=\sigma(\operatorname{Pooling}(\boldsymbol{k}_{t}))$
@@ 3.2 Liger: Linearizing LLMs to Gated Recurrent Structures
Figure 3: Hybrid Architecture of Liger. Liger adopts intra-hybrid Liger Attention and inter-hybrid model architecture by stacking a layer of standard attention Transformer blocks every a few (e.g. 7) layers of Liger Transformer blocks.
@@
LoLCATs-Llama-3-8B 0.04 80.1 80.4 53.5 63.4 72.9 23.0 62.2 70.0
Liger-GLA-Llama-3-8B (Ours) 0.02 80.0 78.7 51.9 76.6 71.4 43.4 67.0 71.7
Table 2: Comparison with Linearized LLMs. Liger outperforms other linearization method on language modeling and understanding tasks with less training tokens across Mistral-7B and Llama-3-8B LLM architectures.
@@ 3.3 Liger Attention: A Hybrid of Sliding Window Attention and Gated Recurrent Modeling
Building upon previous works that combine softmax attention with linear attention (Katharopoulos et al., 2020; MiniMax et al., 2025), we propose a intra-layer hybrid attention mechanism, termed Liger Attention. This method integrates a hybrid form of Gated Recurrent Modeling (GRM) and Sliding Window Attention (SWA) (Beltagy et al., 2020) with narrow softmax attention window size, by blending their outputs in a weighted manner. Specifically, the formulation is given by:
@@
where $\operatorname{GRM}$ denotes Gated Recurrent Modeling (and its variants) in Eq. 3 and $\operatorname{SWA}$ refers to Sliding Window Attention, which is a variant of softmax attention as expressed in Eq. 1 formulated in Eq. 10. The parameters $\alpha$ and $\beta$ control the relative contributions of each attention mechanism.
@@ 3.4 Liger Architecture and Its Hybrid
Liger-GLA-Llama-3-8B (Ours) 0.02 80.0 78.7 51.9 76.6 71.4 43.4 67.0 71.7
Liger-GLA-Llama-3-8B-H (Ours) 0.02 80.4 80.1 52.6 75.8 71.5 44.4 67.5 72.1
Table 3: Performance Comparison of Pre-trained LLM Architectures on Commonsense Reasoning and Knowledge Benchmarks. Results span Transformer-based (Mistral-7B, Llama-3-8B), linear/subquadratic (Mamba, RWKV), hybrid (Zamba), and our linearized Liger-GLA variants on language modeling and understanding tasks. Our Linearized-Liger models achieve competitive performance with only 0.02B training tokens, demonstrating efficient adaptation to gated linear recurrent architectures.
@@ 4.1 Experimental Setups
Liger-GLA-Llama-3-8B (Ours) 67.0 71.7
@@
Liger-GLA $\mathbf{S}_{t}=\mathbf{G}_{t}\odot\mathbf{S}_{t-1}+\boldsymbol{k}_{t}^{\top}\boldsymbol{v}_{t}$ $\boldsymbol{o}_{t}=\boldsymbol{q}_{t}\mathbf{S}_{t}$ $\mathbf{G}_{t}\in\mathbb{R}^{D}$ 71.7 43.4
@@
We conducted experiments to compare efficiency in terms of decoding latency speed and GPU memory consumption of Llama-3-8B without (w/o.) Flash-Attention-2 (FA2), Llama-3-8B with (w/.) Flash-Attention-2 (FA2), Liger-GLA-8B and Liger-GLA-8B-Hybrid architecture (Liger-GLA-8B-H) on single A800 80GB GPU. We set a fixed batch size of 16 and variate the decoding sequence length from 1K to 32K with the fixed prefix input length of 128. As presented in Fig. 4, we observe that Liger achieves linear-time decoding complexity while maintaining constant memory usage.
@@ 4.4 Liger Framework Analysis ($\mathcal{RQ}\mathbf{4}$)
To verify the key components of Liger, we conducted ablation studies on the model structure. Specifically, we experimented with using gates generated from randomly initialized gate projections (Gate Proj.) instead of pooling. Additionally, we incorporated learnable feature map modules (Feat. Map.), similar to Zhang et al. (2024b). We also evaluated the effects of removing LoRA (w/o LoRA) and SWA (w/o SWA) individually. The results of these experiments are detailed in Table 6, demonstrating the effectiveness of proposed components in Liger.
@@ 5 Related Work
Linear Recurrent Models and their Hybrid. To address the challenges of quadratic complexity computation cost in standard softmax attention, many linear recurrent models are proposed to achieve efficient training and inference. Data-dependent gating/decay has been proved as an effective mechanism to control the memory changes and improve sequence modeling expressiveness (Yang et al., 2023; Peng et al., 2024; Qin et al., 2024c; Zhang et al., 2024c; Du et al., 2025). Recently, some layer-wise hybrid model architectures have been proposed to compensate for the lack of memory capacity in the linear recurrent models, such as StripedHyena-Nous (Poli et al., 2023), Jamba (Lieber et al., 2024) and Zamba (Glorioso et al., 2024). In addition, integration of softmax and linear attention shows great potential as a new hybrid attention paradigm, such as Agent Attention (Han et al., 2024), GSA (Zhang et al., 2024c) and Minimax-01 (MiniMax et al., 2025). However, all of these hybrid attention architectures introduce extra modules (e.g. feature mapping and gate modules) that need to be trained from scratch, which increases the complexity of the model architecture design and may lead to suboptimal softmax attention approximation.
@@ Appendix A Datasets and Benchmarks
  • Cleaned Alpaca (Taori et al., 2023): The cleaned Alpaca dataset is a structured dataset designed for instruction-tuning of language models, containing 52,000 unique instructions and corresponding outputs generated by OpenAI’s text-davinci-003 engine. Each entry in the dataset includes an "instruction" field that specifies the task for the model, an optional "input" field providing context or additional information, and an "output" field with the model’s response. The dataset is formatted in JSON and is intended to enhance the ability of language models to follow instructions effectively.
@@ Appendix B Experiment Details
Appendix C Full Results

### v2

@@ Abstract
Transformers with linear recurrent modeling offer linear-time training and constant-memory inference. Despite their demonstrated efficiency and performance, pretraining such non-standard architectures from scratch remains costly and risky. The linearization of large language models (LLMs) transforms pretrained standard models into linear recurrent structures, enabling more efficient deployment. However, current linearization methods typically introduce additional feature map modules that require extensive fine-tuning and overlook the gating mechanisms used in state-of-the-art linear recurrent models. To address these issues, this paper presents Liger, short for Linearizing LLMs to gated recurrent structures. Liger is a novel approach for converting pretrained LLMs into gated linear recurrent models without adding extra parameters. It repurposes the pretrained key matrix weights to construct diverse gating mechanisms, facilitating the formation of various gated recurrent structures while avoiding the need to train additional components from scratch. Using lightweight fine-tuning with Low-Rank Adaptation (LoRA), Liger restores the performance of the linearized gated recurrent models to match that of the original LLMs. Additionally, we introduce Liger Attention, an intra-layer hybrid attention mechanism, which significantly recovers 93% of the Transformer-based LLM at 0.02% pre-training tokens during the linearization process, achieving competitive results across multiple benchmarks, as validated on models ranging from 1B to 8B parameters.
@@ 1 Introduction
Large language models (LLMs) have demonstrated exceptional performance across various natural language processing tasks (Chintala, 2023; Team, 2023; Zhu et al., 2024; Qu et al., 2024). However, the Transformer-based architecture (Vaswani et al., 2017) used in modern LLMs, with its reliance on softmax attention, suffers from quadratic computational complexity. This inefficiency results in significant speed and memory challenges, particularly during pretraining on long sequences. During inference, the Key-Value (KV) cache (Kwon et al., 2023) grows linearly with the input sequence length, leading to reduced inference speed and high memory usage, which severely limits the capability of these models for handling long-sequence tasks (Sun et al., 2024a). In contrast, models based on linear recurrent modeling (Katharopoulos et al., 2020; Yang et al., 2023; Qin et al., 2024b; Sun et al., 2025; Du et al., 2025) provide linear-time training and constant-memory inference, offering substantial efficiency benefits and positioning themselves as promising candidates for the next generation of foundational architectures (MiniMax et al., 2025).
footnotetext: The source code is available at https://github.com/OpenSparseLLMs/Linearization and the models are available at https://huggingface.co/collections/linear-moe-hub.
@@
While pretraining LLMs using architectures based on linear recurrent modeling reduces costs due to their linear training complexity, the high expenses of pretraining from scratch associated with large model sizes and datasets still remain a major obstacle to their adoption and practical use. This challenge has hindered the advancement of linear recurrent models. Linearizing pretrained LLMs like SUPRA (Mercat et al., 2024), MambaInLlama (Wang et al., 2024) and LoLCATs (Zhang et al., 2024a), as an emerging new direction, allows the transfer of weights from an existing pretrained model to one with linear recurrent modeling architectures at a small fraction of the original pretraining cost. The linearization approach is a promising post-training technique to enable efficient pretrained model deployment while preserving their performance. Gating mechanisms (Qin et al., 2024a; Sun et al., 2023) play a crucial role in linear recurrent models by controlling memory retention and forgetting, with their effectiveness widely demonstrated in such architectures. However, incorporating gate modules as additional components requires both transferring weights from pre-trained LLMs and training these gating modules from scratch. This process not only increases the cost of linearization but also creates a larger architectural divergence from Transformer-based LLMs. This divergence may hinder the effective approximation of softmax attention, limiting the performance of gated linear recurrent models (Zhang et al., 2024d). Moreover, existing linearization methods often overlook the detailed design considerations of gated linear models, and the newly added modules fail to leverage the pre-trained weights of LLMs, further reducing the efficiency of linearization.
@@
In this paper, we present Liger, which stands for Linearizing large language models to gated recurrent structures, a novel approach for linearizing LLMs. Liger repurposes the weights from pre-trained Transformer-based LLMs and introduces a novel method for constructing crucial gating mechanisms in gated recurrent structures using the key projection. This approach avoids the complex attention transfer process found in existing linearization methods. After transforming the weights and constructing the gating mechanisms, Liger requires only lightweight fine-tuning of the linearized gated recurrent model parameters through LoRA autoregressive training. By introducing Liger Attention, this efficient process restores further improves the model’s performance with minimal linearization cost, achieving competitive results across a range of language modeling and understanding benchmarks while benefiting from the linear-time inference efficiency of the recurrent architecture.
@@
  • We apply Liger to linearize the latest Llama-3 series, ranging from 1B to 8B parameters. Experimental results show that Liger outperforms existing linearization methods (like SUPRA (Mercat et al., 2024), MambaInLlama (Wang et al., 2024) and LoLCATs (Zhang et al., 2024a)), in terms of both efficiency and its ability to preserve the original performance of pre-trained LLMs.
@@ 2 Preliminary
The standard softmax attention is highly reliant on the growing KV Cache (Chou et al., 2024b; Wang et al., 2024) to recall the history "memory" for sequence modeling, which results in quadratic complexity and costly memory requirements especially in long context setting.
@@
Although linear attention with causal mask matrix cannot use matrix associativity to reduce the parallel form training complexity from quadratic to linear, its chunk-wise parallel form allows hardware-efficient sub-quadratic and partially parallel training (Yang et al., 2023; Sun et al., 2024a, 2025; Qin et al., 2024a).
@@ 3.1 Gate Construction by Key Projection
Gated Retention (Sun et al., 2024b) $\mathbf{G}_{t}=\alpha_{t}\mathbf{1}^{\top}\mathbf{1}$ $\alpha_{t}=\sigma(\operatorname{Pooling}(\boldsymbol{k}_{t}))$
@@ 3.2 Liger: Linearizing LLMs to Gated Recurrent Structures
Figure 3: Liger Hybrid Architecture. Liger adopts intra-hybrid Liger Attention and inter-hybrid model architecture by stacking a layer of standard attention Transformer blocks every a few (e.g. 7) layers of Liger Transformer blocks.
@@
LoLCATs-Llama-3-8B 0.04 80.1 80.4 53.5 63.4 72.9 42.1 65.4 70.0
Liger-GLA-Llama-3-8B (Ours) 0.02 80.3 81.1 52.5 76.3 72.0 43.4 67.6 72.4
Table 2: Linearized LLMs Comparison. Liger outperforms other linearization method on language modeling and understanding tasks with less training tokens across Mistral-7B and Llama-3-8B LLM architectures.
@@ 3.3 Liger Attention: A Hybrid of Sliding Window Attention and Gated Recurrent Modeling
Building upon previous works that combine softmax attention with linear attention (Katharopoulos et al., 2020; Arora et al., 2024; MiniMax et al., 2025), we propose an intra-layer hybrid attention mechanism, termed Liger Attention. This method integrates a hybrid form of Gated Recurrent Modeling (GRM) and Sliding Window Attention (SWA) (Beltagy et al., 2020) with narrow softmax attention window size, by blending their outputs in a weighted manner. Specifically, the formulation is given by:
@@
where $\operatorname{GRM}$ denotes Gated Recurrent Modeling (and its variants) in Eq. 3 and $\operatorname{SWA}$ refers to Sliding Window Attention, which is a variant of softmax attention as expressed in Eq. 1 formulated in Eq. 10. The parameters $\alpha$ and $\beta$ control the relative contributions of each attention mechanism. We found that setting the sum of the two parameters $\alpha$ and $\beta$ to 1 is particularly critical for linearization (simply sets to 0.5 each in our implementation), which can better approximate the original attention output distribution.
@@ 3.4 Liger Architecture and Its Hybrid
Liger-GLA-Llama-3-8B (Ours) 0.02 80.3 81.1 52.5 76.3 72.0 43.4 67.6 72.4
Liger-GLA-Llama-3-8B-H (Ours) 0.02 80.6 80.7 52.7 76.9 71.4 44.4 67.8 72.5
Table 3: Performance Comparison of Pre-trained and Linearized LLMs on Common-sense Reasoning and Knowledge Benchmarks.s Results span Transformer-based (Mistral-7B, Llama-3-8B), linear/subquadratic (Mamba, RWKV), hybrid (Zamba), and our linearized Liger-GLA variants on language modeling and understanding tasks. Our Linearized Liger models achieve competitive performance with only 0.02B training tokens, demonstrating efficient adaptation to gated linear recurrent architectures.
@@ 4.1 Experimental Setups
Liger-GLA-Llama-3-8B (Ours) 67.6 72.4
@@
Liger-GLA $\mathbf{S}_{t}=\mathbf{G}_{t}\odot\mathbf{S}_{t-1}+\boldsymbol{k}_{t}^{\top}\boldsymbol{v}_{t}$ $\boldsymbol{o}_{t}=\boldsymbol{q}_{t}\mathbf{S}_{t}$ $\mathbf{G}_{t}\in\mathbb{R}^{D}$ 72.4 43.4
@@
We conducted experiments to compare efficiency in terms of decoding latency speed and GPU memory consumption of Llama-3-8B without (w/o.) Flash-Attention-2 (FA2), Llama-3-8B with (w/.) Flash-Attention-2 (FA2) and Liger-GLA-8B on single A800 80GB GPU. We set a fixed batch size of 16 and variate the decoding sequence length from 1K to 32K with the fixed prefix input length of 128. As presented in Fig. 4, we observe that Liger achieves linear-time decoding complexity while maintaining constant memory usage.
@@ 4.4 Liger Framework Analysis ($\mathcal{RQ}\mathbf{4}$)
To verify the key components of Liger, we conducted ablation studies on the model structure. Specifically, we experimented with using gates generated from randomly initialized gate projections (Gate Proj.) instead of pooling, and adopting pure linear attention without considering gating mechanism (Pure LA). Additionally, we incorporated learnable feature map modules (Feat. Map.), similar to Zhang et al. (2024b). We also evaluated the effects of removing LoRA (w/o LoRA), GLA (w/o GLA) and SWA (w/o SWA) individually. The results of these experiments are detailed in Table 6, demonstrating the effectiveness of proposed components in Liger.
@@ 5 Related Work
Linear Recurrent Models and their Hybrid. To address the challenges of quadratic complexity computation cost in standard softmax attention, many linear recurrent models are proposed to achieve efficient training and inference. Data-dependent gating/decay has been proved as an effective mechanism to control the memory changes and improve sequence modeling expressiveness (Yang et al., 2023; Peng et al., 2024; Qin et al., 2024c; Zhang et al., 2024c; Du et al., 2025). Recently, some layer-wise hybrid model architectures have been proposed to compensate for the lack of memory capacity in the linear recurrent models, such as StripedHyena-Nous (Poli et al., 2023), Jamba (Lieber et al., 2024), Zamba (Glorioso et al., 2024), Hymba (Dong et al., 2024), Titans (Behrouz et al., 2024) and Minimax-01 (MiniMax et al., 2025). Actually, all of these hybrid attention architectures introduce extra modules (e.g. feature mapping or gate modules) that need to be trained from scratch, which increases the complexity of the model architecture design and may lead to suboptimal softmax attention approximation. In addition, integration of softmax and linear attention shows great potential as a new intra-layer hybrid attention paradigm, such as Agent Attention (Han et al., 2024), Based (Arora et al., 2024), GSA (Zhang et al., 2024c). However, these inter-layer hybrid attention forms have linear recurrent modules that only focus on long-term modeling and ignore local information (Arora et al., 2024), and lack a gating mechanism and the approximation of the attention output distribution by controlling the hybrid ratio (or hybrid weights) (Han et al., 2024; Arora et al., 2024), which is particularly critical in the linearization process.
@@ 6 Conclusion
Acknowledgements
This work is supported by the Shanghai AI Laboratory.
@@ Appendix A Datasets and Benchmarks
  • Cleaned Alpaca (Taori et al., 2023): The cleaned Alpaca dataset is a structured dataset designed for instruction-tuning of language models, containing 52,000 instructions and corresponding outputs generated by OpenAI’s text-davinci-003 engine. Each entry in the dataset includes an "instruction" field that specifies the task for the model, an optional "input" field providing context or additional information, and an "output" field with the model’s response. The dataset is formatted in JSON and is intended to enhance the ability of language models to follow instructions effectively.
@@ Appendix B Experiment Details
LoLCATs Liger
$\parallel\Delta W\parallel_{1}$ 3.68e+06 2.91e+06
${\parallel\Delta W\parallel_{1}}/{W}$ 4.82% 3.85%
$\parallel\Delta W\parallel_{2}$ 3.93e+04 1.27e+03
${\parallel\Delta W\parallel_{2}}/{W}$ 179.29% 7.63%
Table 7: Model Weight Changes after Linearization. Liger has a smaller number of changing model weights than LoLCATs.
Appendix C Model Weight Changes
We analyze the model weight changes after model architecture linearization. Let $W$ be the sum of each weight parameter of LLM, we calculate the model weight changes $\Delta W$ in Liger compared with LoLCATs. As presented in Table 7, we observe that Liger has a smaller number of changing parameters than LoLCATs, which mitigates catastrophic forgetting of pretrained knowledge during linearization process, thereby enhancing both training efficiency and model effectiveness.
Appendix D Gate Construction from Different Components
We consider constructing gating mechanisms from different components of the model. We compare the Liger-GLA-8B’s performance of gate module obtained by pooling from query (Q), key (K) and value (V) matrices, and the results are shown in Table 8. We observed that the gate construction from key matrix outperforms than others. The key matrix usually serves as memory indexing in transformer or linear model (Gershman et al., 2025), which is similar to the gate that determines which part of the memory should be retrieved or forgotten. In this view, it is intuitive to choose Key for gate construction. MetaLA (Chou et al., 2024a) also points out that the linear recurrent model needs to meet the "least parameter approximation" condition to achieve the optimal linear attention design. In this case, the parameters of key matrix are redundant and can be used to construct the gate, which also provides motivation and theoretical support for gate derived from Key.
Model PiQA ARC-e ARC-c Hella. Wino. MMLU Avg. Avg.
Liger-GLA-8B-Q-Pooling 80.2 78.6 51.6 76.7 71.6 41.7 66.7 71.7
Liger-GLA-8B-K-Pooling 80.3 81.1 52.5 76.3 72.0 43.4 67.6 72.4
Liger-GLA-8B-V-Pooling 80.7 77.4 50.6 76.4 70.4 43.7 66.5 71.1
Table 8: Results on Gate Construction from Different Components. We compare the Liger-GLA-8B’s performance of gate module obtained by pooling from query (Q), key (K) and value (V) matrices, the gate construction from key matrix outperforms than others.
Appendix E Full Results